OpenStudy (anonymous):

someone help me What is the volume of a cone with a radius of 6 cm and a height of 12 cm? Use 3.14 for pi. Round your answer to the nearest hundredth. 452.16 cm^3 465.24 cm^3 472.12 cm^3 480.50 cm^3

OpenStudy (anonymous):

Do you know the formula for area of a cone?

OpenStudy (cwrw238):

volume of a cone = (1/3)*pi*r^2*h

OpenStudy (cwrw238):

so plug in your values for radius r and height h and work it out

OpenStudy (anonymous):

V=3.14*6^2*(12/3)= 452.16

OpenStudy (anonymous):

okay I worked my prob but I got another question I need help on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

.What is the volume of a cylinder with a radius of 1.5 inches and a height of 19 inches? Use 3.14 for pi. Round your answer to the nearest hundredth. 28.5 cubic inches 100.25 cubic inches 134.24 cubic inches 156.73 cubic inches